# Amortizing Perpetual Options

### v1

@@ 3 Risk-Neutral Valuation
Table 1: Summary table of AmPO risk-neutral valuation and Greeks for a constant notional exposure.
Within this section, we want to investigate the risk-neutral valuation of these AmPOs as presented in the prior section. To do so, we follow Assumption 3.1 throughout the remainder of this work.
@@
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space satisfying the usual conditions. Consider a complete market with constant risk-free rate $r\geq 0$ and risky asset with price $S=(S_{t})_{t\geq 0}$. Let $\mathbb{Q}\sim\mathbb{P}$ denote the unique risk-neutral measure for this market and consider the $\mathbb{Q}$-Brownian motion (w.r.t. $(\mathcal{F}_{t})_{t\geq 0}$) $W=(W_{t})_{t\geq 0}$. For the remainder of this work, we will assume that $S$ follows a geometric Brownian motion under $\mathbb{Q}$:
@@
with initial price $S_{0}>0$ and volatility $\sigma>0$.
3.1 Pricing
To determine the risk-neutral value of an AmPO, we first observe a fundamental symmetry in the value process. Holding an AmPO involves a decaying notional $N_{t}=e^{-qt}N_{0}$ on the options payoff $\Phi(\cdot)$. That is, from the holder’s perspective, the position’s value is $N_{t}V(S_{t})$ for AmPO option value $V(\cdot)$ with payoff $N_{\tau}\Phi(S_{\tau})$ at exercise time $\tau$. The amortization immediately shows up as an additional discounting term (i.e., an effective risk-free rate $r+q$); with this augmented risk-free rate, the amortization also functionally appears as a drag on the modified drift to the underlying asset (i.e., an effective dividend yield $q$). In this way, in Lemma 3.2, we demonstrate that AmPO valuation perfectly coincides with (vanilla) perpetual American options under modified parameters.
Lemma 3.2.
The AmPO with amortization rate $q>0$ is priced identically to the perpetual American option with the same payoff, effective risk-free rate $r+q$, and dividend yield $q$.
Proof.
Let $V:\mathbb{R}_{+}\to\mathbb{R}$ denote the price of one unit of the AmPO as a function of the underlying. Given an original notional position $N_{0}=1$ wlog, the AmPO contract’s value at time $t$ is given by $\Pi_{t}=e^{-qt}V(S_{t})$. In particular, the discounted value process $(e^{-rt}\Pi_{t})_{t\geq 0}$ must be a $\mathbb{Q}$-martingale. Applying Itô’s lemma, we recover the standard Black-Scholes differential equation:
$0=\frac{1}{2}\sigma^{2}S_{t}^{2}V^{\prime\prime}(S_{t})+rS_{t}V^{\prime}(S_{t})-(r+q)V(S_{t}).$
The desired result follows directly by comparing this ODE to the standard result for perpetual American options on dividend-paying assets (see, e.g., [11, Chapter 26.2]). ∎
Following Lemma 3.2, the pricing of call and put AmPOs trivially follow from the results of, e.g., [11, Chapter 26.2]. For this reason, we omit the proof for the pricing of call and put AmPOs.
@@ 3.1 Pricing
We conclude this discussion of risk-neutral valuation by considering how the amortization $q$ can mimic an effective maturity date $T$. That is, for a given amortization rate, we can define the effective maturity as the time $T$ such that the option’s premium coincides with a standard, dated American option with the same underlying dynamics.
@@
Consider an asset price following Assumption 3.1 with risk-free rate $r=5\%$, volatility $\sigma=50\%$, and initial price $S_{0}=100$. Consider the at-the-money call option (i.e., $K=100$) with amortization rate $q$. In Figure 1(a), we plot the maturity of the at-the-money (American) call option which has the same premium as the AmPO with amortization rate qq. [^1] As the assets under consideration here pay 0 dividends, American and European call options have the same premium. As expected, the effective maturity $T$ is decreasing in the amortization rate $q$. Notably, even though the effective maturity is decreasing, the effective notional ($e^{-qT}$) at the effective maturity is also decreasing in the amortization rate (see Figure 1(b)). For the selected parameters, even at high amortization rates ($q\approx 1$), the AmPO has over 65% of its original notional value remaining at the time that the equivalent dated option would expire.
(a) The effective maturity date $T$ as a function of the amortization rate $q$.
(b) The effective notional ($e^{-qT}$) of the AmPO with amortization rate $q$ at its effective maturity date $T$.
@@
3.2 Greeks
As call and put AmPOs can be valued via an equivalence relation with perpetual American options as provided in Section 3.1, we can follow standard results to determine the dependence of these structures on the underlying system parameters. The main results from this section are summarized in Table 1. Throughout this section we omit the proofs as they follow directly from the aforementioned relation to vanilla perpetual American options.
Remark 4.
Herein we consider the Greeks of the valuation for a constant notional exposure, i.e., dependence of the explicit premia $C_{0}$ and $P_{0}$ for calls and puts, respectively. However, we note that the realized exposure is amortizing for an investor that is either long or short AmPOs. Therefore, in practice, the realized Greeks should be discounted by the amortization factor appropriately. (Further discussions will be provided for Theta in Remark 5 below.)
Corollary 3.6 (Delta).
Consider call and put AmPOs with strike $K>0$ and amortization rate $q>0$. The Delta for call and put options are provided, respectively, by:
$\Delta_{C}:=\frac{\partial C_{0}}{\partial S_{0}}=\begin{cases}\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{\alpha_{C}-1}&\text{if }S_{0}\leq\bar{S}_{C}\\ 1&\text{if }S_{0}>\bar{S}_{C}\end{cases}\quad\text{ and }\quad\Delta_{P}:=\frac{\partial P_{0}}{\partial S_{0}}=\begin{cases}-\left(\frac{\alpha_{P}K}{(1+\alpha_{P})S_{0}}\right)^{1+\alpha_{P}}&\text{if }S_{0}\geq\bar{S}_{P}\\ -1&\text{if }S_{0}<\bar{S}_{P}.\end{cases}$
Furthermore, $\Delta_{C}>0$ and $\Delta_{P}<0$ for any $S_{0}>0$.
Corollary 3.7 (Gamma).
Consider call and put AmPOs with strike $K>0$ and amortization rate $q>0$. The Gamma for call and put options are provided, respectively, by:
$\Gamma_{C}:=\frac{\partial^{2}C_{0}}{\partial S_{0}^{2}}=\begin{cases}\frac{(\alpha_{C}-1)^{2}}{\alpha_{C}K}\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)^{\alpha_{C}-2}&\text{if }S_{0}\leq\bar{S}_{C}\\ 0&\text{if }S_{0}>\bar{S}_{C}\end{cases}\;\text{ and }\;\Gamma_{P}:=\frac{\partial^{2}P_{0}}{\partial S_{0}^{2}}=\begin{cases}\frac{\alpha_{P}K}{S_{0}^{2}}\left(\frac{\alpha_{P}K}{(1+\alpha_{P})S_{0}}\right)^{\alpha_{P}}&\text{if }S_{0}\geq\bar{S}_{P}\\ 0&\text{if }S_{0}<\bar{S}_{P}.\end{cases}$
Furthermore, $\Gamma_{C},\Gamma_{P}\geq 0$ for any $S_{0}>0$.
Remark 5 (Theta).
Theta, typically, refers to the dependence of the premium on the “time to maturity”. Because AmPOs are perpetual options, the risk-neutral price exhibits zero explicit time decay (e.g., $\frac{\partial C_{0}}{\partial t}=0$ for call options). However, due to the amortization, the holder or underwriter of such options will, in fact, have an economic exposure that depends on time. This decay is directly computable as $qC_{0}<0$ or $qP_{0}<0$ for call and put options, respectively. It is exactly this non-zero Theta(-like) decay that distinguishes AmPOs from vanilla perpetual options and creates an effective realized time exposure.
Corollary 3.8 (Vega).
Consider call and put AmPOs with strike $K>0$ and amortization rate $q>0$. The Vega for call and put options are provided, respectively, by:
$\nu_{C}$ $=\frac{\partial C_{0}}{\partial\sigma}=\begin{cases}\frac{4C_{0}}{\sigma}\log\left(\frac{(\alpha_{C}-1)S_{0}}{\alpha_{C}K}\right)\left(\frac{(\alpha_{C}-2)r-q}{(2\alpha_{C}-1)\sigma^{2}+2r}\right)&\text{if }S_{0}\leq\bar{S}_{C}\\ 0&\text{if }S_{0}>\bar{S}_{C}\end{cases}$
$\nu_{P}$ $=\frac{\partial P_{0}}{\partial\sigma}=\begin{cases}\frac{4P_{0}}{\sigma}\log\left(\frac{(1+\alpha_{P})S_{0}}{\alpha_{P}K}\right)\left(\frac{(2+\alpha_{P})r+q}{(2\alpha_{P}+1)\sigma^{2}-2r}\right)&\text{if }S_{0}\geq\bar{S}_{P}\\ 0&\text{if }S_{0}<\bar{S}_{P}.\end{cases}$
Furthermore, $\nu_{C},\nu_{P}\geq 0$ for any $S_{0}>0$.
Example 3.9.
@@ 3.2 Greeks
Consider the same setting as Example 3.5. Again comparing AmPOs to their effective dated counterparts, we want to consider the relative safety of AmPOs. In Figure 2(a), we plot the “safety ratio” between the AmPO Gamma and the Gamma of the effectively dated (American) call option. The lower this ratio, the more stable the option price is to fluctuations in the underlying. Notably, this ratio is increasing in the amortization rate, but stays below 80% for even extremely high amortization rates (e.g., $q=1$). Similarly, in Figure 2(b), we consider the “cost efficiency” of AmPOs by computing the ratio between the economic Theta (i.e., $qC_{0}$) from amortization against the Theta of the effectively dated option. Here, again, we find that the dated options lose value faster even at high amortization rates with a maximum observed efficiency ratio (at $q=1$) of under 75%.
(a) The ratio of the AmPO Gamma against its effectively dated peer.
(b) The ratio of the AmPO (economic) Theta against its effectively dated peer.
3.3 Comparative Statistics on Amortization Rate
@@
Finally, as the amortization rate $q$ appears directly within both the effective risk free rate ($r+q$) and effective dividend yield ($q$) of the equivalently valued perpetual American option, the choice of this contractual parameter can have a strong influence on the valuation and Greeks of AmPOs. However, since we have proposed AmPOs to maintain a constant $q$ throughout the lifetime of the option, we do not want to consider this dependence in the same way as the Greeks (i.e., for hedging purposes), but rather as a way to understand the implications and tradeoffs of choosing the $q$ for a contract. In doing so, we will focus on three aspects of the AmPO contracts: the risk-neutral valuation; the optimal exercise boundary; and the Vega. Throughout this section we set $\bar{\alpha}:=(\alpha_{C}+\alpha_{P})/2>0$ for notational simplicity.
@@
First, as expected, increasing $q$ accelerates economic amortization, reducing the value for the holder and, therefore, also the premium. This is provided in Corollary 3.11 below.
@@
In the prior Examples 3.5 and 3.9, we have analyzed how AmPOs compare with dated options. In the following example, we wish to consider how selecting the amortization rate can be used to optimize a volatility-based strategy.
@@
Consider the setting of Example 3.5. Recall from Corollary 3.11 and Corollary 3.13 that both the premium and Vega of an AmPO are decreasing in the amortization rate $q$. As the realized Vega of a position depends on the budget actually invested, we now want to consider how the ratio of the Vega to premium behaves to consider a simplified optimization problem: consider an investor who wants to take as leveraged a volatility position as possible using AmPOs with a \$100 budget constraint. To do so, we consider three strategies: bullish volatility with call options (Figure 3(a)); bearish volatility with put options (Figure 3(b)); and neutral volatility with an at-the-money straddle (Figure 3(c)). Notably, though bullish and neutral strategies have increasing positional Vegas, the bearish strategy has a non-monotonic behavior coming from the simultaneous drop in both premium and Vega. For the current parameter setting, the optimal bearish volatility strategy (which also provides the maximum positional Vega among all 3 strategies) is with an amortization rate of $q^{*}\approx 14.26\%$.
@@ 4 Discussion and Applications
Thus far, we have considered the construction and valuation of AmPOs. Within this section, we will briefly consider use cases for these novel perpetual options. In doing so, we will consider applications both within traditional financial markets and decentralized, i.e., blockchain-based, markets.
4.1 Traditional Finance
@@ 5 Conclusion
Within this work, we introduced amortizing perpetual options, a fungible variant of continuous-installment options which are, therefore, suitable for exchange-based trading. Under a constant amortization rate, we found an equivalence relation between AmPOs and vanilla perpetual American options on dividend-paying assets. This relation permits standard analytical valuations and, thus, allowed us to characterize the Greeks and analyze the sensitivities of option behavior to the contractual amortization rate.
Our analysis in Section 3 assumed the complete market structure of the Black-Scholes framework with the constant amortization rate. Extending the study of AmPOs to incomplete markets and stochastic amortization rates presents several natural directions for future work. For instance, one could investigate the optimal amortization rate for long and short positions as there is a natural tradeoff between a higher amortization rate and lower initial premium (Corollary 3.11). In this way, an equilibrium amortization rate can arise from utility-indifference or heterogeneous-agent considerations. Such extensions would provide a framework for designing amortization rates in practice.

### v2

@@ 3.2 Constant Amortization in the Black-Scholes Framework
Table 1: Summary table of AmPO risk-neutral valuation and Greeks for a constant notional exposure in the Black-Scholes framework under constant amortization.
Within this section, we want to investigate the risk-neutral valuation of these AmPOs in the Black-Scholes framework under constant amortization. To do so, we follow Assumptions 3.4 throughout this subsection.
@@
Consider a complete market with constant risk-free rate $r\geq 0$ and risky asset with price $S=(S_{t})_{t\geq 0}$. Let $\mathbb{Q}\sim\mathbb{P}$ denote the unique risk-neutral measure for this market and consider the $\mathbb{Q}$-Brownian motion (w.r.t. $(\mathcal{F}_{t})_{t\geq 0}$) $W=(W_{t})_{t\geq 0}$. For the remainder of this work, we will assume that $S$ follows a geometric Brownian motion under $\mathbb{Q}$:
@@
with initial price $S_{0}>0$ and volatility $\sigma>0$. Furthermore, consider AmPOs with constant amortization rates $q_{t}\equiv q>0$.
3.2.1 Pricing and Hedging
Following Lemma 3.2, the pricing of call and put AmPOs trivially follow from the results of, e.g., [16, Chapter 26.2]. For this reason, we omit the proof for the pricing of call and put AmPOs.
@@ 3.2.1 Pricing and Hedging
As call and put AmPOs can be valued via an equivalence relation with perpetual American options as provided in Lemma 3.2 and the above explicit representations, we can follow standard results to determine the dependence of these structures on the underlying system parameters. Such results are summarized in Table 1.
Remark 3.
Via the relation to perpetual American options, we consider the Greeks of the valuation for a constant notional exposure, i.e., dependence of the explicit premia $C_{0}$ and $P_{0}$ for calls and puts, respectively. However, we note that the realized exposure is amortizing for an investor that is either long or short AmPOs. Therefore, in practice, the realized Greeks should be discounted by the amortization factor appropriately.
This is especially prominent for Theta which, typically, refers to the dependence of the premium on the “time to maturity”. Because AmPOs are perpetual options, the risk-neutral price exhibits zero explicit time decay (e.g., $\frac{\partial C_{0}}{\partial t}=0$ for call options). However, due to the amortization, the holder or underwriter of such options will, in fact, have an economic exposure that depends on time. This decay is directly computable as $qC_{0}<0$ or $qP_{0}<0$ for call and put options, respectively. It is exactly this non-zero Theta(-like) decay that distinguishes AmPOs from vanilla perpetual options and creates an effective realized time exposure.
We conclude this discussion of risk-neutral valuation by considering how the amortization $q$ can mimic an effective maturity date $T_{\text{eff}}$. That is, for a given amortization rate $q$, we define the effective maturity $T_{\text{eff}}:=\inf\{T\geq 0\;|\;V_{0}(S_{0};r,\sigma,q)\leq V^{American}(S_{0};T,r,\sigma)\}$ as the maturity $T$ for a dated American option with the same market dynamics ($S_{0},r,\sigma$) which matches the AmPO valuation. [^1] The effective maturity $T_{\text{eff}}\in(0,\infty)$ for any $q\in(0,\infty)$ is such that $V_{0}(S_{0};r,\sigma,q)=V^{American}(S_{0};T_{\text{eff}},r,\sigma)$ as $V^{American}(S_{0};0,r,\sigma)=\Phi(S_{0})<V_{0}(S_{0};r,\sigma,q)<\lim_{T\to\infty}V^{American}(S_{0};T,r,\sigma)$ by Lemma 3.2; see also Remark 4.
@@
(a) The effective maturity date $T_{\text{eff}}$ as a function of the amortization rate $q$.
(b) The effective notional ($e^{-qT_{\text{eff}}}$) of the AmPO with amortization rate $q$ at its effective maturity date $T_{\text{eff}}$.
(c) The ratio of the AmPO Gamma against its effectively dated peer.
(d) The ratio of the AmPO (economic) Theta against its effectively dated peer.
@@
Consider an asset price following Assumption 3.4 with risk-free rate $r=5\%$, volatility $\sigma=50\%$, and initial price $S_{0}=100$. Consider the at-the-money call option (i.e., $K=100$) with amortization rate $q$.
In Figure 1(a), we plot the effective maturity $T_{\text{eff}}$ of the at-the-money (American) call option which has the same premium as the AmPO with amortization rate qq. [^2] As the assets under consideration here pay 0 dividends, American and European call options have the same premium. As expected, the effective maturity $T_{\text{eff}}$ is decreasing in the amortization rate $q$. However, Figure 1(b) reveals that $T_{\text{eff}}$ decreases at a slower rate than $q$ increases. Consequently, the effective notional (the claimable notional $e^{-qT_{\text{eff}}}$ remaining at the effective maturity $T_{\text{eff}}$) is decreasing in the amortization rate $q$. Crucially, this effective notional illustrates that this effective maturity is solely a pricing equivalence and does not represent even an approximate termination of the AmPO; unlike the dated option which expires at $T_{\text{eff}}$, the AmPO retains a substantial portion of its notional at that time. For the selected parameters, even at high amortization rates ($q\approx 1$), the AmPO has over 65% of its original notional value remaining at the time that the equivalent dated option would expire.
@@
Though the premiums of the AmPO and its effectively dated counterpart coincide, their sensitivities do not because the underlying stopping problems are distinct. For this reason, we now compare (select) Greeks of AmPOs against their effectively dated counterparts. First, in Figure 1(c), we plot the ratio of the AmPO Gamma to the Gamma of the effectively dated (American) call option, i.e., $\Gamma_{0}(S_{0};r,\sigma,q)/\Gamma^{American}(S_{0};T_{\text{eff}},r,\sigma)$. The lower this ratio, the more stable the option price is to fluctuations in the underlying; for this reason we call this the “safety ratio.” Notably, this safety ratio is increasing in the amortization rate, but stays below 80% for even extremely high amortization rates (e.g., $q\approx 1$). Similarly, in Figure 1(d), we consider the “cost efficiency” of AmPOs by computing the ratio between the economic Theta (i.e., $qC_{0}$) from amortization against the Theta of the effectively dated option, i.e., $qC_{0}/\Theta^{American}(S_{0};T_{\text{eff}},r,\sigma)$. For a cost efficiency of less than 1, the AmPO’s loss in notional from amortization is less than the loss in the effectively dated option from the passage of time. Here, again, we find that the dated options lose value faster even at high amortization rates with a maximum observed efficiency ratio (at $q=1$) of under 75%.
3.2.2 Comparative Statics on Amortization Rate
@@
As the amortization rate $q$ appears directly within both the effective risk free rate ($r+q$) and effective dividend yield ($q$) of the equivalently valued perpetual American option, the choice of this contractual parameter can have a strong influence on the valuation and Greeks of AmPOs. However, since we have proposed AmPOs to maintain a constant $q$ throughout the lifetime of the option, we do not want to consider this dependence in the same way as the Greeks (i.e., for hedging purposes), but rather as a way to understand the implications and tradeoffs of choosing the $q$ for a contract. In doing so, we will focus on three aspects of the AmPO contracts: the risk-neutral valuation; the optimal exercise boundary; and the Vega. Throughout this section we set $\bar{\alpha}:=(\alpha_{C}+\alpha_{P})/2>0$ for notational simplicity.
@@
First, as follows from [22, Proposition 5] when applied to (3.1) directly, increasing $q$ accelerates economic amortization, reducing the value for the holder and, therefore, also the premium. This is proven explicitly in the Black-Scholes setting of 3.4 below.
@@
Remark 5.
Though we have provided Corollaries 3.9 and 3.10 for the Black-Scholes setting only, an application of [22, Proposition 5] indicate that the monotonicity of the dependence on the amortization rate $q$ found in those results hold also for general price processes.
@@
In Example 3.7, we analyzed how AmPOs compare with dated options. In the following example, we wish to consider how selecting the amortization rate can be used to optimize a volatility-based strategy.
@@
Consider the setting of Example 3.7. Recall from Corollary 3.9 and Corollary 3.11 that both the premium and Vega of an AmPO are decreasing in the amortization rate $q$. As the realized Vega of a position depends on the budget actually invested, we now want to consider how the ratio of the Vega to premium ($\nu_{0}/V_{0}$) behaves to consider a simplified optimization problem: consider an investor who wants to take as leveraged a volatility position as possible using AmPOs with a \$100 budget constraint. To do so, we consider three strategies: bullish volatility with at-the-money call options (Figure 2(a)); bearish volatility with at-the-money put options (Figure 2(b)); and neutral volatility with an at-the-money straddle (Figure 2(c)). First, note that while the Vega is decreasing in the amortization rate (Corollary 3.11) in all three strategies, the cost of a constant notional position is also decreasing in the amortization rate (Corollary 3.9). In this way, as observed in Figure 2, the total positional Vega depends on the interaction between the Vega $\nu_{0}$ and the option premium $V_{0}$. Notably, though bullish and neutral strategies have increasing positional Vegas, the bearish strategy has a non-monotonic behavior coming from the simultaneous drop in both premium and Vega. For the current parameter setting, the optimal bearish volatility strategy (which also provides the maximum positional Vega among all 3 strategies) is with an amortization rate of $q^{*}\approx 14.26\%$.
@@ 4 Discussion and Applications
4.1 Setting the Amortization Rate
In Section 3.2, we imposed a fixed $q_{t}\equiv q>0$ amortization rate. While the general theory permits endogenous amortization (i.e., dependent on the option value), such a construction can introduce novel manipulations in AmPOs in practice. For instance, consider an AmPO with the fixed dollar installment rate $c_{t}\equiv c$, i.e., $q_{t}=c/V_{t}$, as is often taken for CI options (see, e.g., Example 2.1). For a deep out-of-the-money AmPO ($V_{t}\searrow 0$), the amortization rate diverges to infinity ($q_{t}\nearrow\infty$) causing the open interest in the option series to crash to 0. This is the economically justified result if the fair value is near 0; however, if the price is temporarily manipulated towards the 0 lower bound then this attack can, effectively, wipe out all long option positions. In particular, within decentralized finance where price manipulations are a constant threat, this endogenous amortization rate is particularly fragile. In contrast, the fixed amortization rate $q_{t}\equiv q$ is not subject to such manipulations as the amortization schedule is invariant to market prices. As such, though only a special case of AmPOs (Definition 2.2), we recommend setting the amortization rate as a fixed constant in practice.
Once the design space is restricted to the fixed amortization rate AmPOs, it still remains necessary to choose this rate $q>0$. As AmPOs are exchange-tradable instruments, it is necessary to impose standard contracts in order to concentrate liquidity in specific markets; this is akin to setting a standard choice of maturity date for options (e.g., the third Friday of the month). As evidenced by Example 3.12, it is clear that there does not exist a Pareto optimal amortization rate $q$ amongst all market participants. Instead, herein we suggest the market operator define a few canonical rates (e.g., 3 bps/day, 10 bps/day, and 50 bps/day) to segment market participants based on their implied maturities and risk tolerances (see, e.g., Figure 1). A formal study of optimal choice of amortization rates to maximize, e.g., total market volume when underwriters and option holders compete in an equilibrium setting goes beyond the scope of this work.
4.2 Applications
Thus far, we have considered the construction and valuation of AmPOs. Within this section, we will briefly consider use cases for these novel perpetual options. In doing so, we will consider applications both within traditional financial markets and decentralized, i.e., blockchain-based, markets. In line with the discussion in Section 4.1 above, we restrict our discussion herein to the constant amortization rate setting of Section 3.2.
4.2.1 Traditional Finance
@@ 5 Conclusion
Within this work, we introduced amortizing perpetual options, a fungible variant of continuous-installment options which are, therefore, suitable for exchange-based trading. Notably, we found an equivalence relation between AmPOs and vanilla perpetual American options on dividend-paying assets. This relation permits standard analytical valuations and, thus under constant amortization rates, allowed us to characterize the Greeks and analyze the sensitivities of option behavior to the contractual amortization rate.
Our analysis assumed pre-determined amortization rate functions. Future work could investigate the optimal amortization rate for long and short positions as there is a natural tradeoff between a higher amortization rate and lower initial premium (Corollary 3.9). In this way, an equilibrium amortization rate can arise from utility-indifference or heterogeneous-agent considerations. Such extensions would provide a framework for designing amortization rates in practice that go beyond the discussion in Section 4.1.